a wheel turning with angular velocity of 60π rad/s is brought to rest with a constant acceleration. it makes 60 rotations before it stops .calculate its angular velocity.
1
Expert's answer
2020-11-16T07:45:40-0500

If the wheel makes 60 rotations before it stops then it coveres angular distance


"\varphi = 60\times 2\pi =120\space rad"

before it stops. On the other hand, the angular distance covered in time "t" is:


"\varphi =\omega_0t -\dfrac{\varepsilon t^2}{2}"

where "\omega_0 = 60\pi \space rad/s" is the initial angular velocity and "\varepsilon" is the constant angular acceleration. Expressing "\varepsilon", obtain:


"\varepsilon = \dfrac{2}{t^2}(\omega_0t-\varphi)"

The angular velocity in any given moment of time is:


"\omega(t) = \omega_0-\varepsilon t = \omega_0 - \dfrac{2}{t}(\omega_0t-\varphi) = \dfrac{2\varphi}{t} - \omega_0"

Substituting values, obtain:

"\omega(t) = \dfrac{2\varphi}{t} - \omega_0 = \dfrac{2\cdot 120\pi}{t} - 60\pi = 240\pi\dfrac{1}{t} - 60\pi\space [rad/s]"

Answer. Angular velocity at any given moment of time (in seconds) is "\omega(t) = 240\pi\dfrac{1}{t} - 60\pi\space [rad/s]"